You are the campaign manager for a prominent politician. It is 8 weeks before the election and you would like to make some predictions on how the voters will vote. You go out and get a random sample of 800 register voters that are likely to vote in the upcoming election and find out that 240 of them will vote for your candidate. Calculate a 99% confidence interval to estimate the true proportion that are likely to vote for your candidate.

a) (28.32% , 31.68%)

b) (27.33% , 32.67%)

c) (26.82% , 33.18%)

d) (25.83% , 34.17%)
Answer #1

solution : Ộ=2 240 n = 8001 2=246, = 0.3 Given - n 800 (=99.1 -0.99 x=1-2 -1-0.99 =0.01 10.0 -2.576 At 991 C1=-{{+ 2 x SP (1-